OpenStudy (anonymous):

Solve the system of equations. Use the substitution method to solve. 4x – 2y = 24 x – y = 3

OpenStudy (anonymous):

take -2y in right hand side

OpenStudy (anonymous):

4x – 2y = 24 x – y = 3 ----> x = 3+y = y+3 4x – 2y = 24 4(y+3) – 2y = 24 4y+12 – 2y = 24 2y+12 = 24 2y= 12 y=6 can you solve for x ?

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

You welcome !

OpenStudy (texaschic101):

if it makes it any easier for you, the first equation..4x - 2y = 24 can reduce to 2x - y = 12. It will probably make it a little easier using smaller numbers

OpenStudy (anonymous):

Ok thank you.

OpenStudy (texaschic101):

sure thing :)